OpenStudy (anonymous):

supply missing term to make expression a perfect trinomial. 9f^2- ___+25

OpenStudy (anonymous):

30x

OpenStudy (anonymous):

How do you do this?

OpenStudy (anonymous):

notice that 9f^2 = (3f)^2 and 25 = 5^2 from the formula (a-b)^2 we know that (a-b)^2 = a^2 - 2ab + b^2

OpenStudy (anonymous):

im still not 100% on it.

OpenStudy (anonymous):

9f^2- ___+25 = (3f)^2 - ____+5^2 to make this into a perfect trinomial, you need to complete the square. which is what I did using the formula (a-b)^2 where a = 3f and b = 5

OpenStudy (anonymous):

so you subtract 9f^2 and -4 and get 5 and add 25. Wouldnt it be 30f^2?

OpenStudy (anonymous):

what is (3f-5)^2? expand it and post the answer.

OpenStudy (anonymous):

-2f^2=-4f

OpenStudy (anonymous):

what? do you know the formula for (a-b)^2